OpenStudy (bbb911):

Find the ratio of the width of a regulation size basketball court to its perimeter. 50 ft (width) 94 ft (length)

jimthompson5910 (jim_thompson5910):

P = 2(L+W) P = 2(94+50) P = ???

OpenStudy (bbb911):

i was wondering where did you get the 2?

jimthompson5910 (jim_thompson5910):

well to find the perimeter of any rectangle, you add the 4 sides BUT the opposite sides are equal

jimthompson5910 (jim_thompson5910):

so you could do it the long way 94+94+50+50

jimthompson5910 (jim_thompson5910):

or you could write it as 2*94 + 2*50

jimthompson5910 (jim_thompson5910):

then factor out the 2 2*(94 + 50)

OpenStudy (bbb911):

yea the perimeter is 288 right cos i added all the sides but im stuck right there

jimthompson5910 (jim_thompson5910):

now compute the ratio width : perimeter 50 : 288 25 : 144 ... divide both parts by 2

OpenStudy (bbb911):

yea like i dont know were the 2 comes from is it automatically there?

jimthompson5910 (jim_thompson5910):

do you see how I went from 94+94+50+50 to 2*94 + 2*50

OpenStudy (bbb911):

where i am stuck at is getting the 2 like i had no idea that you had to include the 2 that is why my solution was wrong

jimthompson5910 (jim_thompson5910):

you got 50:288 as your answer right?

OpenStudy (bbb911):

well what i did was divide 50 and 288 and i got 5 38/50 and my solution was wrong. i was wondering how 2 is involved like what you put 2 times 94 and 2 times 50

jimthompson5910 (jim_thompson5910):

all they want is the ratio of width:perimeter

jimthompson5910 (jim_thompson5910):

so you replace width with 50 and perimeter with 288 to get 50:288

jimthompson5910 (jim_thompson5910):

then you divide both parts by 2 to get 25:144

jimthompson5910 (jim_thompson5910):

and that's as far as you can go

OpenStudy (bbb911):

okay so for every question like this would i have to divide by 2 ?

jimthompson5910 (jim_thompson5910):

I only divided by 2 because 50 and 288 have the common factor of 2

jimthompson5910 (jim_thompson5910):

The idea was to reduce the ratio as much as possible

jimthompson5910 (jim_thompson5910):

25 and 144 have no factors in common (other than 1), so that's why 25:144 is fully reduced now

OpenStudy (bbb911):

ohhhhhhhh! know i get it i did the total opposite thank you so much ! :)

jimthompson5910 (jim_thompson5910):

you're welcome glad it's clicking now

OpenStudy (bbb911):

yea it did now thanks so much

jimthompson5910 (jim_thompson5910):

sure thing